OpenStudy (anonymous):

(3,5) and (9,9) slope intercept

OpenStudy (anonymous):

first i know that you do (9-5)/(9-3)= 4/6 that is reduced to 2/3

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

Thats the slope u have calculated

OpenStudy (anonymous):

but it says its not right...so my answer is not finished..

OpenStudy (anonymous):

Now, use this formula (y-y1)=m(x-x1)

OpenStudy (anonymous):

Can u? (x1,y1) is the coordiante of the point..... U can use anyone

OpenStudy (anonymous):

And remember slope intercept form must be in y=mx+c

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

im stuck at the last answer i gave you :-(